OpenStudy (anonymous):

Imagine that you are writing a note to your best friend to explain how to add and subtract integers. Write your note, assuming that your friend really needs to know the sign rules for adding and subtracting integers and that he/she must rely on you and only you for an explanation.

OpenStudy (anonymous):

Imagine a number line with a zero in the middle, negative numbers to the left and positive numbers to the right: _|_____|_____|____|____|____|___|____|___|___|____|___ -5 -4 -3 -2 -1 0 1 2 3 4 5 If you add a positive, your starting number will increase to the right 5+2=7 -5+2=-3 If you add a negative your starting number will decrease to left 5-2=3 -5-2=-7 If you subtract a positive your starting number will decrease to left 5-2=3 -5-2=-7 IF you subtract a negative your starting number will increase to the right 5+2=7 -5+2=-3 Notice that subtracting a negative is the same as adding a positive, and Adding a negative is the same as subtracting a positive